# Symplectic mapping class relations from pencil pairs

### v2

@@ 3.1. Pencil pairs
Let $X^{\pm}$ be closed varieties equipped with very ample line bundles $\operatorname{\mathcal{L}}^{\pm}$, Lefschetz pencils $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$ for the $\operatorname{\mathcal{L}}^{\pm}$, and associated subvarieties $B^{\pm}\subset Z^{\pm}_{i}\subset X^{\pm}$. If there is a symplectomorphism
@@
It is implicit in our notation that the pencil pair depends only on the $\operatorname{\mathcal{L}}^{\pm}$ rather than the specified sections $(\boldsymbol{\mathfrak{s}}_{0}^{\pm},\boldsymbol{\mathfrak{s}}_{1}^{\pm})$. This abuse of notation is permitted by Lemma 2.4.1.
@@ 3.2. The case $\dim_{\mathbb{C}}X^{\pm}=2$
Suppose that $X$ is a closed variety of $\dim_{\mathbb{C}}X=2$ equipped with a very ample line bundle $\operatorname{\mathcal{L}}$ having a pencil with associated holomorphic map $\pi$. Then our divisor $Z_{1}$ is a Riemann surface and base locus $B\subset Z_{1}$ is a set of points. Applying Poincaré duality and the adjunction formula,
(3.2.1) #$\begin{gathered}\#B=[Z_{1}]^{2}=\int_{Z_{1}}\omega_{\operatorname{\mathcal{L}}}=\int_{X}c_{1}(\operatorname{\mathcal{L}})^{2},\\ \chi(Z_{1})=\int_{Z_{1}}c_{1}(X)-[Z_{1}]^{2}=\int_{X}(c_{1}(X)-c_{1}(\operatorname{\mathcal{L}}))\wedge c_{1}(\operatorname{\mathcal{L}})\\ \implies\#\operatorname{Crit}\pi=\int_{X}c_{2}(X)+(c_{1}(\operatorname{\mathcal{L}})-2c_{1}(X))\wedge c_{1}(\operatorname{\mathcal{L}}).\end{gathered}$
Since the symplectomorphism type of $(Z_{1},\omega_{\operatorname{\mathcal{L}}},B)$ is then determined by $\chi(Z_{1})$, the symplectic area of $Z_{1}$, and #$\#B$ (which is redundant), the following theorem is obvious.
@@
Let $X^{\pm}$ be a pair of $\dim_{\mathbb{C}}=2$ varieties equipped with very ample line bundles ±$\operatorname{\mathcal{L}}^{\pm}\to X^{\pm}$. Then there is a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair iff the quantities
@@ 3.3. Generalized lantern relations
    $Z_{0,\chi,d,k}=\boldsymbol{\mathfrak{s}}_{0}^{-1}(0)$ with $[Z_{0,\chi,d,k}]=dk^{2}$ and $\chi(Z_{0,\chi,d,k})=k(\chi+d-dk)$,
@@
    #$\#\operatorname{Crit}=3dk^{2}+2(\chi-kd-k\chi)$ critical points of $\pi=\boldsymbol{\mathfrak{s}}_{1}/\boldsymbol{\mathfrak{s}}_{0}:X_{\chi,d,k}\to\mathbb{C}$, and
@@ 4.1. Projective models
Suppose that the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have divisors with a common projective model. Then there are pencils for the ±$\operatorname{\mathcal{L}}^{\pm}\rightarrow X^{\pm}$ and a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
@@
By the definition of projective models there are sections $\boldsymbol{\mathfrak{s}}^{\pm}_{\mathcal{M},1}$ of $\mathcal{O}_{\vec{d}}$ with zero loci $Z_{\mathcal{M}}^{\pm}\subset\mathcal{M}$ and biholomorphisms $\phi^{\pm}:Z^{\pm}_{1}\rightarrow Z_{\mathcal{M}}^{\pm}$ for which ±$\operatorname{\mathcal{L}}^{\pm}|_{Z_{1}^{\pm}}=\phi^{\ast}\mathcal{O}_{d_{1}}$. Choose sections $\boldsymbol{\mathfrak{s}}^{\pm}_{0}$ of the $\operatorname{\mathcal{L}}^{\pm}$ so that the $(\boldsymbol{\mathfrak{s}}^{\pm}_{0},\boldsymbol{\mathfrak{s}}^{\pm}_{1})$ are pencils for the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ with base loci
@@ 4.2. Cabling pencil pairs
Let $Z\subset X$ be a divisor of be a very ample line bundle $\operatorname{\mathcal{L}}$ over a closed variety $X$. View $Z$ and $X$ as symplectic manifolds with their symplectic structures inherited from $(\mathbb{CP}^{N},\omega_{\mathbb{CP}^{N}}),N=\dim H^{0}(X,\operatorname{\mathcal{L}})-1$ via the embedding of Equation (2.4.1). For $k\geq 0$, let $Z_{k}$ be a divisor of $\operatorname{\mathcal{L}}^{\otimes k}=\mathcal{O}_{k}|_{X}$ intersecting $Z$ transversely. Then as a symplectic manifold, $Z_{k}\subset X\subset\mathbb{CP}^{N}$ is a $k$-fold branched covering of $Z$ with branch locus $Z\cap Z_{k}$. In particular, the symplectomorphic type of the pairs $(Z_{k},Z\cap Z_{k})$ and $(Z_{k},B_{k})$, where $B_{k}$ is the base locus of a pencil on $X$ containing $Z_{k}$, depends only on $\operatorname{\mathcal{L}}|_{Z}$.
@@
This is a $k$-fold branched covering whose branch locus is the zero section. If $T_{k}$ is the tautological section of $p_{k}^{\ast}\mathcal{O}_{k}$ over $\operatorname{Tot}(\mathcal{O}_{k}|_{Z})$ where $p_{k}:\operatorname{Tot}(\mathcal{O}_{k}|_{Z})\to Z$ is the projection, then $\operatorname{pow}^{k}Z^{\prime}_{k}$ is the zero locus of $\boldsymbol{\mathfrak{s}}_{k}+T_{k}$. This zero locus is exactly the graph of $-\boldsymbol{\mathfrak{s}}_{k}$ in $\operatorname{Tot}(\mathcal{O}_{k}|_{Z})$. Hence the branch locus of $\operatorname{pow}_{k}|_{Z^{\prime}_{k}}$ is exactly $\{\boldsymbol{\mathfrak{s}}_{k}=0\}\cap Z=Z^{\prime}_{k}\cap Z$. Therefore $p_{k}\circ\operatorname{pow}_{k}:Z^{\prime}_{k}\to Z$ is a holomorphic branched covering with branch locus $Z_{k}^{\prime}\cap Z=Z_{k}\cap Z$ as claimed. ∎
@@
This follows immediately from Theorem 4.2.1 and the definitions involved. The projective models for polarized divisors $(Z^{\pm}_{1},\operatorname{\mathcal{L}}^{\pm})$ are deformation equivalent through complete intersection subvarieties of some $\mathcal{M}\subset\mathbb{CP}^{N}$ for which $\operatorname{\mathcal{L}}^{\pm}$ are pull-backs of the line bundle $\mathcal{O}_{d_{1}}\to\mathbb{CP}^{N}$. Theorem 4.2.1 then provides that for all $k$, pairs of divisors and branch-loci for $(X^{\pm},(\operatorname{\mathcal{L}}^{\pm})^{\otimes k})$ will be symplectomorphic by some $\phi_{k}$. Such $\phi_{k}$ define the desired pencil pairs. ∎
@@ 5. $\dim_{\mathbb{C}}X^{\pm}=3$ examples
A simply connected $\dim_{\mathbb{C}}=2$ Calabi-Yau variety is a $K3$ surface. According to Kodaira, the diffeomorphism type of a $K3$ surface is uniquely determined. See [H16] for this and additional background. For an ample line bundle $\operatorname{\mathcal{L}}$ over a $K3$ surface $Z$, the genus is defined as
@@
which computes the genus of a divisor in $Z$ cut out by a generic holomorphic section of $\operatorname{\mathcal{L}}$. It follows from a Hodge theory computation that $\chi(Z)=24$. The adjunction formula applied to anticanonical divisors for a $\dim_{\mathbb{C}}X=3$ Fano variety then gives
@@
If $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a pencil associated to a line bundle $\operatorname{\mathcal{L}}$ over a variety $X$ of $\dim_{\mathbb{C}}X=3$, then the associated Lefschetz fibration on $X\setminus\{\boldsymbol{\mathfrak{s}}_{0}=0\}$ has
#$\#\operatorname{Crit}\pi=-\chi(X)+\int_{X}2c_{2}(X)c_{1}(\operatorname{\mathcal{L}})-3c_{1}(X)c_{1}(\operatorname{\mathcal{L}})^{2}+4c_{1}(\operatorname{\mathcal{L}})^{3}$
@@
To count #$\#\operatorname{Crit}\pi$ we seek to apply the formula from Theorem 2.4.2, which requires computations of $\chi(Z_{1})$ and $\chi(B)$ where $Z_{1}$ is a divisor of $X$ cut out by some generic section $\boldsymbol{\mathfrak{s}}_{1}$ of $\operatorname{\mathcal{L}}\to X$ and $B$ is the base locus of a pencil. Throughout we apply axiomatic properties of Chern classes, cf. [MiS74].
Since the $\operatorname{\mathcal{L}}$-valued $1$-form $d\boldsymbol{\mathfrak{s}}_{1}$ is non-degenerate along $Z_{1}$, the coimage of the map $TX|_{Z_{1}}\to\operatorname{\mathcal{L}}|_{Z_{1}}$, $v_{z}\mapsto d\boldsymbol{\mathfrak{s}}_{1}(v_{z})$ is a normal bundle to $Z_{1}\subset X$. Thus we have an identification $TX|_{Z_{1}}\simeq TZ_{1}\oplus\operatorname{\mathcal{L}}|_{Z_{1}}$ which when combined with the multiplicativity of total Chern classes yields the adjunction formula
@@
The first equality follows from the fact that $c_{2}(Z_{1})$ is the Euler class of $Z_{1}$. The third equality follows from $[Z_{1}]=\operatorname{PD}_{X}c_{1}(\operatorname{\mathcal{L}})$. Similar arguments provide us with
@@
Finally we apply $-\operatorname{Crit}\pi=\chi(X)-2\chi(Z_{1})+\chi(B)$ from Theorem 2.4.2 and add up the computed terms to obtain the general formula. In the Fano case with $\operatorname{\mathcal{L}}=A_{X}^{\otimes k}$, Equation (5.0.1) yields
@@ 5.2. Examples of pencil pairs from anticanonical divisors on Fano $3$-folds
Let $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ be polarized Fano $3$-folds for which $A_{X^{\pm}}=(\operatorname{\mathcal{L}}^{\pm})^{\operatorname{ind}_{X^{\pm}}}$ and $\operatorname{\mathcal{L}}^{\pm}$ are very ample. Suppose that their anticanonical divisors $Z^{\pm}\subset X^{\pm}$ are such that $g(Z^{\pm},\operatorname{\mathcal{L}}^{\pm})\leq 10$, or equivalently
@@ 5.2.6. Case $\operatorname{ind}_{X}=2,\int_{X}c_{1}(A_{X})^{3}=48\implies g(Z,\operatorname{\mathcal{L}})=7$
    $X$ is a divisor of $\deg=(1,1)$ in $\mathbb{CP}^{2}\times\mathbb{CP}^{2}$ with $\chi(X)=6$. Here $\operatorname{\mathcal{L}}$ is of degree $(1,1)$ so that $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Since $\operatorname{\mathcal{L}}$ is the pullback of $\mathcal{O}_{1}$ from the Segre embedding $\mathbb{CP}^{2}\times\mathbb{CP}^{2}\rightarrow\mathbb{CP}^{8}$ and so is very ample.
@@
    $X=(\mathbb{CP}^{1})^{3}$ having $\chi(X)=8$. Here $\operatorname{\mathcal{L}}$ has $\deg=(1,1,1)$ giving $\operatorname{\mathcal{L}}^{\otimes 2}=A_{X}$. Again $\operatorname{\mathcal{L}}$ is very ample as it is determined by a Segre embedding.
@@
The complement of a divisor for $\operatorname{\mathcal{L}}$ in the $X\subset\mathbb{CP}^{2}\times\mathbb{CP}^{2}$ indexed (2-32) can be identified with a cotangent disk bundle $\mathbb{D}^{\ast}\mathbb{CP}^{2}$, and is studied from a symplectic point of view in [Au07, Bi01, T20]. In particular, [T20, §5] describes a Lefschetz fibration associated to $\operatorname{\mathcal{L}}$ on this Weinstein $8$-manifold.
@@ 5.3. Oba’s example revisited
As in §5.2.6 above, consider $(X^{+},\operatorname{\mathcal{L}}^{+})$ given by a $\deg=(1,1)$ hypersurface in $\mathbb{CP}^{2}\times\mathbb{CP}^{2}$ equipped with $\operatorname{\mathcal{L}}^{+}$ a $\deg=(1,1)$ line bundle and $(X^{-},\operatorname{\mathcal{L}}^{-})$ is $(\mathbb{CP}^{1})^{3}$ with $\operatorname{\mathcal{L}}^{-}$ a $\deg=(1,1,1)$ line bundle. Then divisors $Z^{\pm}_{1}\subset X^{\pm}$ for the line bundles ±$\operatorname{\mathcal{L}}^{\pm}\to X^{\pm}$ are del Pezzo surfaces (that is, Fano $2$-folds) of degree $\int_{Z_{1}^{\pm}}A_{Z_{1}^{\pm}}^{2}=6$. In these cases the ±$\operatorname{\mathcal{L}}^{\pm}|_{Z_{1}^{\pm}}$ are the anticanonical bundles $A_{Z^{\pm}_{1}}$. Smoothly, we have equivalences
@@
The number of critical points of a Lefschetz fibration $\pi^{-}_{l}$ associated to $(X^{-},(\operatorname{\mathcal{L}}^{-})^{\otimes l})$ may then be computed using the first formula in Lemma 5.0.1. From Theorem 2.4.2 we can count the number of critical points of a Lefschetz fibration $\pi^{+}_{l}$ associated $(X^{+},(\operatorname{\mathcal{L}}^{+})^{\otimes l})$ using #$\#\operatorname{Crit}\pi^{-}_{l}$ and the $\chi(X^{\pm})$. The counts are
#$\#\operatorname{Crit}\pi^{\pm}_{l}=-\chi(X^{\pm})+12l(2l^{2}-3l+2),\quad\chi(X^{+})=6,\quad\chi(X^{-})=8.$

### v3

@@ 3.1. Pencil pairs
[…]
@@
[…]
@@ 3.2. The case $\dim_{\mathbb{C}}X^{\pm}=2$
[…]
(3.2.1) $\begin{gathered}\#B=[Z_{1}]^{2}=\int_{Z_{1}}\omega_{\operatorname{\mathcal{L}}}=\int_{X}c_{1}(\operatorname{\mathcal{L}})^{2},\\ \chi(Z_{1})=\int_{Z_{1}}c_{1}(X)-[Z_{1}]^{2}=\int_{X}(c_{1}(X)-c_{1}(\operatorname{\mathcal{L}}))\wedge c_{1}(\operatorname{\mathcal{L}})\\ \implies\#\operatorname{Crit}\pi=\int_{X}c_{2}(X)+(c_{1}(\operatorname{\mathcal{L}})-2c_{1}(X))\wedge c_{1}(\operatorname{\mathcal{L}}).\end{gathered}$
Since the symplectomorphism type of $(Z_{1},\omega_{\operatorname{\mathcal{L}}},B)$ is then determined by $\chi(Z_{1})$, the symplectic area of $Z_{1}$, and #B\#B (which is redundant), the following theorem is obvious.
@@
Let $X^{\pm}$ be a pair of $\dim_{\mathbb{C}}=2$ varieties equipped with very ample line bundles $\operatorname{\mathcal{L}}^{\pm}\to X^{\pm}$. Then there is a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair iff the quantities
@@ 3.3. Generalized lantern relations
    $Z_{0,\chi,d,k}=\boldsymbol{\mathfrak{s}}_{0}^{-1}(0)$ with $Z_{0,\chi,d,k}]=dk^{2}$ and $\chi(Z_{0,\chi,d,k})=k(\chi+d-dk)$,
@@
    $\operatorname{Crit}=3dk^{2}+2(\chi-kd-k\chi)$ critical points of $\pi=\boldsymbol{\mathfrak{s}}_{1}/\boldsymbol{\mathfrak{s}}_{0}:X_{\chi,d,k}\to\mathbb{C}$, and
@@ 4.1. Projective models
Suppose that the $(X^{\pm},\operatorname{\mathcal{L}}^{\pm})$ have divisors with a common projective model. Then there are pencils for the $\operatorname{\mathcal{L}}^{\pm}\rightarrow X^{\pm}$ and a $\phi$ for which $(X^{\pm},\operatorname{\mathcal{L}}^{\pm},\phi)$ is a pencil pair.
@@
[…]
@@ 4.2. Cabling pencil pairs
[…]
@@
This is a $k$-fold branched covering whose branch locus is the zero section. If $T_{k}$ is the tautological section of $p_{k}^{\ast}\mathcal{O}_{k}$ over $\operatorname{Tot}(\mathcal{O}_{k}|_{Z})$ where $p_{k}:\operatorname{Tot}(\mathcal{O}_{k}|_{Z})\to Z$ is the projection, then $\operatorname{pow}^{k}Z^{\prime}_{k}$ is the zero locus of $\boldsymbol{\mathfrak{s}}_{k}+T_{k}$. This zero locus is exactly the graph of $\boldsymbol{\mathfrak{s}}_{k}$ in $\operatorname{Tot}(\mathcal{O}_{k}|_{Z})$. Hence the branch locus of $\operatorname{pow}_{k}|_{Z^{\prime}_{k}}$ is exactly $\{\boldsymbol{\mathfrak{s}}_{k}=0\}\cap Z=Z^{\prime}_{k}\cap Z$. Therefore $p_{k}\circ\operatorname{pow}_{k}:Z^{\prime}_{k}\to Z$ is a holomorphic branched covering with branch locus $Z_{k}^{\prime}\cap Z=Z_{k}\cap Z$ as claimed. ∎
@@
[…]
@@ 5. $\dim_{\mathbb{C}}X^{\pm}=3$ examples
[…]
@@
[…]
@@
[…]
$\operatorname{Crit}\pi=-\chi(X)+\int_{X}2c_{2}(X)c_{1}(\operatorname{\mathcal{L}})-3c_{1}(X)c_{1}(\operatorname{\mathcal{L}})^{2}+4c_{1}(\operatorname{\mathcal{L}})^{3}$
@@
To count $\operatorname{Crit}\pi$ we seek to apply the formula from Theorem 2.4.2, which requires computations of $\chi(Z_{1})$ and $\chi(B)$ where $Z_{1}$ is a divisor of $X$ cut out by some generic section $\boldsymbol{\mathfrak{s}}_{1}$ of $\operatorname{\mathcal{L}}\to X$ and $B$ is the base locus of a pencil. Throughout we apply axiomatic properties of Chern classes, cf. [MiS74].
[…]
@@
The first equality follows from the fact that $c_{2}(Z_{1})$ is the Euler class of $Z_{1}$. The third equality follows from $Z_{1}]=\operatorname{PD}_{X}c_{1}(\operatorname{\mathcal{L}})$. Similar arguments provide us with
@@
Finally we apply $\operatorname{Crit}\pi=\chi(X)-2\chi(Z_{1})+\chi(B)$ from Theorem 2.4.2 and add up the computed terms to obtain the general formula. In the Fano case with $\operatorname{\mathcal{L}}=A_{X}^{\otimes k}$, Equation (5.0.1) yields
@@ 5.2. Examples of pencil pairs from anticanonical divisors on Fano $3$-folds
[…]
@@ 5.2.6. Case $\operatorname{ind}_{X}=2,\int_{X}c_{1}(A_{X})^{3}=48\implies g(Z,\operatorname{\mathcal{L}})=7$
[…]
@@
[…]
@@
[…]
@@ 5.3. Oba’s example revisited
[…]
@@
The number of critical points of a Lefschetz fibration $\pi^{-}_{l}$ associated to $(X^{-},(\operatorname{\mathcal{L}}^{-})^{\otimes l})$ may then be computed using the first formula in Lemma 5.0.1. From Theorem 2.4.2 we can count the number of critical points of a Lefschetz fibration $\pi^{+}_{l}$ associated $(X^{+},(\operatorname{\mathcal{L}}^{+})^{\otimes l})$ using $\operatorname{Crit}\pi^{-}_{l}$ and the $\chi(X^{\pm})$. The counts are
$\operatorname{Crit}\pi^{\pm}_{l}=-\chi(X^{\pm})+12l(2l^{2}-3l+2),\quad\chi(X^{+})=6,\quad\chi(X^{-})=8.$